Recent Maintenance – MBAR

MBAR operational status:

2024.04.11

Data from the STS-6 is non-seismic as of day 2024,102 (April 11).

2021.12.08

Station operators are onsite and upgraded the VSAT, installed new VSAT modem and repointed antenna, worked on the power system, replaced the WXT restoring the weather station channels, and updated the FW for the spare Q330.

2020.10.24

The weather station channels started gapping excessively on day 2020,298 (10/24).  IDA engineers are working with the station operator to resolve this issue.

2020.02.12

The station operator visited the station and performed maintenance on the power system, including replacing one of the backup batteries, the sequential power distribution box, and power supplies.  The station operator also replaced both station computers and installed a Paroscientific microbarometer.

2019.11.13

The station is back online after the station operator visited the station and restored power.   A faulty Wilmore DC/DC converter was replaced with an Axiomatic version.

2019.10.19

The station is down due to power issues.  IDA engineers are working with the station operator to restore power.

2019.09.24

IDA engineers are onsite and installed a new STS-6 seismometer and Hyperion Infrasound sensor.  They are also upgrading the power system.

2019.08.11

The station is back online as of day 2019,223-0523.  Outages are still expected due to power issues at the station.

2019.08.05

The station is down as of day 2019,216-10:28 due to power issues yet to be determined.  IDA engineers are planning a station visit to upgrade the power system.

2019.07.31

IDA engineers visited the station to replace the KS54000 seismometer with an STS-6 seismometer.  There were problems removing the KS54000 seismometer which prevented installation of the STS-6. Additional equipment is being shipped to the station for STS-6 installation during an upcoming trip to upgrade the power system.

2019.05.19

The station was down from day 2019,110 to 2019,188 due to problems with the station computer.  Data acquisition was restored after the station operator visited the station and rebooted the station computer.  An engineer also restored telemetry by repointing the satellite antenna and upgrading firmware for the satellite modem.

2019.04.20

Telemetry to the station is down since day 2019,110 due to a satellite change for the network.  An engineer is planning to visit the station to re-point the satellite antenna.

2018.10.30

The station was down from day 2018,299-04:53 to 2018,302-21:32 due to loss of the network connection to the station computer.  The station operator visited the station and cycled power to the station computer restoring data acquisition.

2017.12.19

The station is up as of day 2017,352-21:14 (Dec. 18) after the station operator replaced the station computers.

2017.12.05

The station is down as of day 2017,333-11:05 due to failure of the station computers.  New station computers are being prepared to ship to the station.

2016.02.04

A December 2015 maintenance visit to MBAR was made to investigate the loss of the station’s satellite link and carry out repairs necessary to return it to operational status. On-site testing indicated that the Block Up-Converter (BUC), satellite router, and Low-Noise Block Down-Converter (LNB) had failed.  These items were replaced, a dirty feed-horn cleaned, and its cover replaced.  Repairs to Inter-Facility Link (IFC) cable conduit were carried out and new cables laid.

2014.08.26

The station is back online as of day 2014,238-11:35 after the station operator visited the station and restored power.

2014.08.23

The station is down as of day 2014,235-06:34 due to power problems. Lightning damage is suspected, and the station operator is planning to visit the station to investigate.

2014.08.05

IDA engineers are on-site to upgrade the solar power system and replace the seismometers.  New sensors include new KS54000 and Trillium 120PH seismometers, and an episensor strong motion sensor.

2013.04.01

Telemetry to the station has been restored. The E-W channels from the CMG-3T were intermittently non-seismic as a result of power outages since March 13. The sensor was unlocked and successfully re-centered after these outages ceased. The data is seismic as of day 2013,084-21:51.

2013.03.19

Telemetry to the station remains down. In addition to the power problem, there is no communication with the VSAT modem after the VSAT was re-positioned.

2013.03.13

The station operator reports there is a power outage at the station. The station may still be functioning. However, the outage is causing a communications problem between equipment. There are plans for a station visit to fix the power next week.

2012.09.18

Data quality for the East-West channels from the CMG-3T seismometer has improved since the masses for the sensor were re-centered on day 2012,262 (Sept. 18,2012).

2012.09.04

Data from the strong motion sensor at MBAR is non-seismic.

2012.08.26

The East-West channels from the CMG-3T seismometer are showing high background noise since day 2012,239-21:00.

2012.04.19

IDA station MBAR has been upgraded to a Quanterra Q330HR system. Instruments include KS54000 and CMG-3T seismometers, and a strong motion sensor.

2012.01.23

IDA station MBAR is back on-line as of day 2012,023-08:07 after the station operator replaced the DAS and Ethernet extender.

2011.11.17

The station is down as of day 2011,321-11:22 due to problems with the DAS.

2011.10.31

The station is back online as of day 2011,304-11:31. Details concerning the DAS problems at the station are still under investigation.

2011.10.11

The station is down as of day 2011,284-13:52 due to problems with the DAS. A new digitizer and ethernet extenders are being prepared to ship to the station.

2011.10.10

The station is back on-line as of day 2011,283-15:43 after the station operator replaced the DAS. The station operator reports there are still digitizer problems and also a problem with the Iboot.

2011.09.19

The station remains down after the DAS was manually power cycled. A new DAS will be prepared to ship to the station.

2011.08.31

The station is down as of day 2011,243-19:06 due to communication problems between the DAS and the ISP. The station operator has been requested to visit the station to investigate and cycle power to the DAS.

2011.08.09

The station is back on-line as of day 2011,221-12:15 after the station operator cycled power to the DAS. An Iboot has been shipped to the station. Upon installation, power to the DAS can be cycled remotely.

2011.08.03

The station is down as of day 2011,215-08:00 due to communication problems between the DAS and the ISP that are yet to be determined.

2011.07.05

The station is back on-line as of day 2011,186 after the station operator visited the station and cycled power to the digitizer. The North-South channels for the CMG-3T are also back in normal operating range after the station operator re-centered the masses for the CMG-3T.

2011.06.21

The station is down as of day 2011,172-20:20 due to problems with the DAS that are yet to be determined.

2011.06.14

The station is back on-line as of day 2011,165-07:19 after the station operator cycled power to the DAS.

2011.05.24

The station is down as of day 2011,144-17:08 due to problems with the DAS that are yet to be determined.

2011.05.22

The station is back on-line as of day 2011,142-07:01, after the station operator replaced the ISP and network switch given to him on his visit to UCSD. The North-South channels from the CMG-3T are flat-lined, and the wdo channel is currently not being acquired for reasons yet to be determined.

2011.04.04

The station operator reports visiting the station. Telemetry to the station is apparently up, however, the ISP is still unreachable. Further investigation is underway.

2011.03.12

The station is currently unreachable since day 2011,072. IDA engineers have requested that the station operator investigate.

2010.11.04

The VSAT link has been restored after the station operator reconfigured the VSAT modem.

2010.11.01

The station operator replaced the VSAT modem at the station. The VSAT link is still down. Telemetry remains via gci link.

2008.06.30

The station is back online after an IDA engineer visited the station and restored power, replaced the ISP, and restored communication between the DAS and the ISP.

2008.03.25

The station is down due to power problems. A visit by an IDA engineer is planned to restore power to the station.

2008.02.09

The station was down from about day 2008,038 to 2008,040 due to a power problem. Normal operation resumed after a reboot of the DAS.

2008.01.30

The station is back online as of about day 2008,030 after an IDA engineer restored communication between the DAS and the ISP.

2008.01.23

The station is down again as of about day 2008,023 due to communication problems between the DAS and the ISP. An IDA engineer is enroute to the station.

2008.01.19

The station is back online as of about day 2008,019-14:05 after the station operator cycled power to the DAS.

2007.11.30

The station is down as of about day 2007,334 due to communication problems between the DAS and the ISP that are yet to be determined.

2007.10.24

The station is back online after an IDA engineer visited the station and replaced the DAS, and restored power to the VSAT system.

2007.07.20

The station is down as of about day 201-16:07 due to communication problems between the ISP and the DAS that are yet to be determined.

2007.06.16

The DC-DC converter in the CMG-3T seismometer has been replaced and the short period channels are seismic as of about day 167-12:43.

2007.06.13

The station is back up after faulty solar batteries for the power system were replaced.

2007.06.11

The station is down due to problems with the power system that are yet to be determined.

2007.05.23

The short period channels for the CMG-3T seismometer are nonseismic as of about day 143-18:21 for reasons yet to be determined.

2007.03.13

IDA engineers visited the station and replaced the KS-54000 and CMG-3T seismometers, and the FBA-23 strong motion sensor. All channels were restored except the vertical channels from the KS-54000 seismometer due to problems with the sensor possibly caused during shipping.

2007.02.03

The station was down from day 033-08:15 to 034-09:34 due to problems with the ISP. Normal data acquisition resumed after the station computer was rebooted.

2006.08.24

Data quality for all channels from the FBA-23 sensor is poor for reasons yet to be determined.

2006.02.05

The station was down from day 30 (Jan.30) to day 35 (Feb.5) due to problems with the digitizer. The station is back up after the station operator removed faulty circuit boards in the DAS.

2006.01.29

Data quality for the CMG3-T sensor’s East-West channel is poor due to a faulty sensor.

2005.09.22

The broadband channels for the KS-54000 seismometer are out as of 261-01:45 (Sept.18) for reasons yet to be determined.

2004.07.24

An IDA engineer visited the station to correct the channel swapping problem and replace its GPS clock.

2004.06.16

Telemetry data showed GPS clock had problem since 2004:167-18:12. Station operator has been notified.

2004.05.28

Two of the broadband channels, BBE and BBN, have been swapped since 2004139. This problem has been corrected during processing stage. Telemetry broadband BBE and BBN data prior to 2004148-19:00 were having incorrected channels assigned.

2004.05.18

Station is up after an IDA engineer visited the site to replace the damaged Vsat equipment, solar batteries and failed 24 bit A/D boards. Telemetry is also up after the satellite link is repaired.

2003.09.08

Station down. Station operator reports no data from DAS and no clock (9/3). Telemetry still down as well.

2003.07.07

Telemetry down. There are numerous gaps in NRTS data since 2003.06.19.

2003.06.20

The station operator performed the procedure to center the CMG3-T seismometer masses, but there is no indication on the CMG3-T control box that the masses are actually recentered. The horizontal channels remain out of range.

2003.06.11

The CMG-3T seismometer horizontal channels are dead, bh2 data are clipped around 2003:152-08:00, bh1 channel on the NRTS loop is dead around 2003:157-16:19.

2003.05.12

Field engineer replaced solar panels, updated EPROMS, restored telemetry. A fence was erected around the site to increase security. Station operational as of 2003.05.12.

2002.11.15

The operator phoned to report that thieves had stolen 12 solar panels from the station’s electrical power system. It is not known if these were taken from the telemetry subsystem, the elements powering IDA data acquisition equipment, or those powering the JPL/UNAVCO data acquisition system. IDA is attempting to obtain more details.

2002.09.04

The satellite telemetry circuit failed on this date. IDA engineers are attempting to identify which parts have failed and develop a strategy to restore data flow.

2002.08.19

The GPS clock was replaced and is operating correctly.

2002.07.10

East-West channels from the KS-54000 are down as of 2002:102-09:00 (April 12, 2002) along with the vertical channels.

2002.06.13

The GPS clock is now not working properly.

2002.04.22

Station is operating without input from the KS-54000 vertical channels.